# MgNO: Efficient Parameterization of Linear Operators via Multigrid

### v1

@@ C.2 Navier Stokes
  • Original training setup: Samples consist of 20 sequential time steps, with the goal of predicting the subsequent 10 time steps from the preceding 10. Using the roll-out prediction approach as described by [23], the neural operator uses the initial 10-time steps to forecast the immediate next time step. This predicted time step is then merged with the prior 9 time steps to predict the ensuing time step. This iterative process continues to forecast the remaining 10 time steps.
  • New training setup: Our findings indicate that an amalgamation of deep learning strategies is pivotal for the optimal performance of time-dependent tasks. While [23] employed the previous 10-time steps as inputs for the neural operator, our approach simplifies this. We find that leveraging just the current step’s data, similar to traditional numerical solvers, is sufficient. During training, we avoid model unrolling. Instead, from the initial 1000 samples consisting of 20 sequential time steps, we derive 19,000 samples with pairs of sequential time steps. These are then shuffled and used to train the neural operator to predict the subsequent time step based on the current one. For testing, we revert to the roll-out prediction method as only the initial time step’s ground truth is available. These methods align with some techniques presented in [1]. As shown in Table 2, the second approach provides better performance.
@@
This dataset in [22] is devoted to simulate incompressible flow through a pipe. The governing equations are:
@@
C.3 Helmholtz equation
Helmholtz equation in highly heterogeneous media is an example of multiscale wave phenomena, whose solution is considerably expensive for complicated and large geological models. We adopt the setup from [5], for the Helmholtz equation on the domain $D=[0,1]^{2}$. Given frequency $\omega=10^{3}$ and wave speed field $c:\Omega\rightarrow\mathbb{R}$, the excitation field $u:\Omega\rightarrow\mathbb{R}$ solves the equation
@@ C.3 Helmholtz equation
Appendix D Discretization invariance
@@
Appendix E Model Configurations

### v2

@@ D.2 Navier Stokes
  • Original training setup: Samples consist of 20 sequential time steps, with the goal of predicting the subsequent 10 time steps from the preceding 10. Using the roll-out prediction approach as described by Li et al. (2020), the neural operator uses the initial 10-time steps to forecast the immediate next time step. This predicted time step is then merged with the prior 9 time steps to predict the ensuing time step. This iterative process continues to forecast the remaining 10 time steps.
  • New training setup: Our findings indicate that an amalgamation of deep learning strategies is pivotal for the optimal performance of time-dependent tasks. While Li et al. (2020) employed the previous 10-time steps as inputs for the neural operator, our approach simplifies this. We find that leveraging just the current step’s data, similar to traditional numerical solvers, is sufficient. During training, we avoid model unrolling. Instead, from the initial 1000 samples consisting of 20 sequential time steps, we derive 19,000 samples with pairs of sequential time steps. These are then shuffled and used to train the neural operator to predict the subsequent time step based on the current one. For testing, we revert to the roll-out prediction method as only the initial time step’s ground truth is available. These methods align with some techniques presented in Brandstetter et al. (2022). As shown in Table 2, the second approach provides better performance.
@@
This dataset in Li et al. (2022) is devoted to simulate incompressible flow through a pipe. The governing equations are:
@@
D.3 Helmholtz equation
Helmholtz equation in highly heterogeneous media is an example of multiscale wave phenomena, whose solution is considerably expensive for complicated and large geological models. We adopt the setup from (De Hoop et al., 2022), for the Helmholtz equation on the domain $D=[0,1]^{2}$. Given frequency $\omega=10^{3}$ and wave speed field $c:\Omega\rightarrow\mathbb{R}$, the excitation field $u:\Omega\rightarrow\mathbb{R}$ solves the equation
@@ D.3 Helmholtz equation
Appendix E Discretization invariance
@@
Appendix F Model Configurations